# Melting curve of correlated iron at Earth’s core conditions from machine‑learned DFT+DMFT

### v2

@@ Abstract
Reliable constraints on iron’s melting curve at Earth’s inner‑core boundary require accurate finite‑temperature electronic correlations, yet DFT+DMFT calculations remain too costly for large‑scale thermodynamic sampling. Here, we develop a machine‑learning accelerator for charge self-consistent DFT+DMFT by training E(3)-equivariant graph neural networks to predict the local self-energy and Fermi level from atomic environments, providing an efficient warm start to the DMFT self-consistency loop. Using high‑throughput data for Fe, FeO, and NiO, we obtain a 2–4 times reuduction in DMFT iterations. Leveraging this improvement, we generate correlated energies and forces for Fe at core pressures, train a neural‑network interatomic potential, and determine the melting curve via two‑phase coexistence simulations. We obtain a predicted melting temperature of 6225 K at 330 GPa.
@@
Correlated electron systems occupy a central place in modern condensed-matter physics. Their rich behaviors, including metal–insulator transitions [mitAbrahams1996, mitMasatoshi1998], unconventional magnetism [ucmBroholm2020, ucmLee2013], and high-temperature superconductivity [htsDagotto1994, htsKeimer2015, htsStewart2017], arise from strong local interactions, particularly in transition-metal compounds where lattice, spin, orbital, and charge degrees of freedom are tightly intertwined [tmoAnisimov1997, tmoTokura2000]. Understanding how this interplay governs electronic properties across pressures, temperatures, and compositions remains one of the major challenges in computational materials physics and is essential for problems spanning functional oxides to materials under extreme pressure and temperature conditions.
@@
Fe at the conditions of Earth’s core exemplifies the challenges of simulating correlated systems [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015]. The melting curve of Fe at hundreds of gigapascals controls the temperature at the inner-core boundary (ICB) and thus constrains the thermal history and dynamics of our planet, yet experimental determinations at these pressures are extremely challenging and theoretical predictions span a broad range of temperatures. While recent experiments have begun to converge on values near 6200 K at ICB pressure [emcAnzellini2013, emcKraus2022, emcBalugani2024], theoretical predictions from density functional theory (DFT) calculations remain scattered, with discrepancies often exceeding 1000 K [aimcAlfe2002, aimcAlfe2009, aimcBelonoshko2021, aimcBelonosko2000, aimcGonzalez2023, aimcSola2009, aimcStixrude2014, aimcSun2022, aimcSun2023, aimcWu2024]. A fundamental limitation of standard DFT approaches is their inadequate treatment of dynamical correlations in Fe [ieccPourovskii2017, ieccPourovskii2019, ieccVekilova2015]. Under core conditions, hexagonal close-packed (hcp) Fe and body centered cubic (bcc) Fe retain their Fermi-liquid character while fcc Fe displays markedly non-Fermi liquid behavior, with relative free energy differences changing significantly with temperature [ieccPourovskii2019]. Incorporation of these correlation-induced differences are neglected with conventional exchange-correlation functionals, but must be included for accurate simulations of liquid phases.
@@
Dynamical mean-field theory (DMFT), which captures local correlations in a fully dynamical and non-perturbative manner [dmftKotliar2006], offers a path forward. Coupled to DFT, it allows for realistic simulations of correlated materials by taking into account effects such as quasiparticle renormalizations, and provides access to electronic and thermodynamic properties such as spectral functions and finite-temperature phase diagrams [dmftArpita2019]. However, this power comes at a substantial computational cost. Each DMFT calculation requires repeated solution of the quantum impurity problem until impurity and lattice occupations become consistent. Determining a melting curve requires sampling thousands of large-scale atomic configurations through molecular dynamics, a task that has remained prohibitively expensive within DFT+DMFT.
@@
Recently, machine learning has begun to be explored as a way to reduce the cost of DMFT calculations, for example through surrogate impurity solvers, learned corrections to approximate Green’s functions, and data-driven models trained on simplified parameter spaces [mlgfMitra2025, mlgfDong2024, mlgfSturm2021, mlgfArsenault2014]. While these studies demonstrate that key DMFT quantities can often be approximated efficiently in selected settings, most existing approaches have been confined to model Hamiltonians or narrow material classes, and the learned targets are not always constructed to respect the analytic structure and symmetry constraints required in realistic DFT+DMFT workflows. It therefore remains unclear whether machine learning can provide a robust, physics-consistent acceleration of DFT+DMFT for correlated solids with the accuracy needed for thermodynamic predictions, particularly under the extreme conditions relevant to Earth’s core.
@@
In this work, we demonstrate that embedding physical constraints into machine-learned self-energies enables, for the first time, systematic DFT+DMFT determination of the iron melting curve at core conditions. Our approach exploits the analytic structure of the self-energy: we decompose it into a static high-frequency limit $\Sigma(\infty)$ and a frequency dependent term $\Sigma(i\omega)$, represented compactly through Legendre coefficients, with each component predicted by E(3)-equivariant graph neural networks that respect the point-group symmetry of the local atomic environment. This physics-informed representation reduces the learning problem to smooth, real-valued targets while preserving the causal structure essential for DMFT self-consistency [mlgfArsenault2014].
@@
The resulting model accelerates DMFT convergence by a factor of two to four across Fe, FeO, and NiO, enabling high-throughput generation of correlated electronic structure data. We leverage this capability to train a machine learned interatomic potential (MLIP) for iron and perform two-phase coexistence simulations spanning 307–352 GPa. Our predicted melting temperature of 6225K at 330 GPa agrees well with the recent experimental measurements [emcAnzellini2013, emcKraus2022, emcBalugani2024], providing strong validation that large-scale molecular dynamics (MD) simulations using DMFT-based MLIPs are viable for phase-equilibrium calculations in planetary interiors.
@@
A naive strategy would be to train a neural network to output $\Sigma(i\omega_{n})$ directly on a discrete Matsubara frequency grid. This is undesirable for several reasons. First, the DMFT self-energy is complex-valued and, when obtained from quantum Monte Carlo impurity solvers, can be noisy at high frequencies. Second, $\Sigma(i\omega_{n})$ obeys strict analytic constraints: it is a causal function with a well-defined high-frequency expansion, $\Sigma(i\omega_{n})=\Sigma(\infty)+{\cal O}(1/i\omega_{n})$ [sigmaWang2011], and its imaginary part has definite sign on the imaginary axis. Direct regression on a large set of complex values risks violating these constraints unless they are enforced explicitly.
@@
$Re\left\{\Sigma(i\omega_{n})\right\}\xrightarrow{\omega_{n}\to\infty}\Sigma(\infty),$ (1)
@@
The dynamical part $\Delta\Sigma(i\omega)$ is then Fourier transformed to imaginary time and expanded in a Legendre polynomial basis [Boehnke2011],
@@
where $\Delta\Sigma^{\mathrm{ML}}(i\omega_{n})$ is obtained by transforming the Legendre expansion back to the Matsubara axis. Both forward and reverse transformations are performed by the TRIQS software package [Parcollet2015, Boehnke2011]. This $\Sigma^{\mathrm{ML}}$, together with the predicted $E_{f}$, provides the initial condition for the DMFT self-consistency loop.
Including $E_{f}$ as a learned quantity is an essential for achieving substantial speedups. In DFT+DMFT the Fermi level is normally adjusted iteratively so that the total electron number satisfies the charge constraint, and poor initial guesses can severely delay convergence. By treating $E_{f}$ on the same footing as $\Sigma(\infty)$ and the Legendre coefficients, we define a joint functional $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ that maps a crystal structure to a complete, physics-consistent set of input parameters.
To implement this functional, we represent each crystal structure as a periodic atomic graph and approximate $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ with an $E(3)$-equivariant graph neural network built using the open-source e3nn software package [e3nnGieger2022, e3nnKondor2018, e3nnThomas2018, e3nnWeiler2018]. Atoms are nodes, connected to neighbors within a radial cutoff, and are encoded by learned embeddings of their chemical species and mass. Relative position vectors between neighbors are expanded in spherical harmonics and radial basis functions, so that all edge features and intermediate representations transform as irreducible representations of $E(3)$. A small stack of equivariant message-passing layers then builds a symmetry-respecting description of each local environment that is translationally invariant and equivariant under rotations and inversion. Further architectural details are given in the Supplemental Material.
@@
To assess the accuracy of the physics-informed functional, we first compare its predictions for the self-energy against fully converged DFT+DMFT results on held-out test structures. For Fe and FeO, structural configurations were generated from MD simulations of high‑symmetry crystalline prototypes (e.g., hcp/bcc/fcc for Fe, B2 phase for FeO) to ensure coverage of both ordered and disordered environments. 2.5‑ps DFT‑MD trajectories (0.5 fs timestep) of 4–8 atom cells were generated and representative snapshots were selected by k‑medoids clustering in structural fingerprint space [fpZhu2016], yielding 500 configurations in total (100-200 per trajectory). Since the fingerprint represents the chemical environment of each crystal structure, the fingerprint distance allows us to quantify the overall difference in chemical environments between any two crystal structures. This ensures that we obtain diverse chemical environments for the solver to “learn” in an attempt to aid generalization of results to unseen structures. For NiO we use configurations from our previous study [Rao2024]. Full details of dataset generation and selection are provided in the Supplemental Material. We reserve 10% as a held-out test set.
@@
Figure 1 (d-f) shows the convergence of impurity-lattice occupation difference $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|$ as a function of DMFT iteration for Fe, FeO, and NiO, averaged over the data for 20 randomly chosen structures. With zero initialization, Fe requires approximately 9 iterations to reach our convergence criterion of $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|<0.05$; with ML-predicted self-energies, convergence is achieved in 2–3 iterations. Similar speedups are observed for FeO (9 vs. 5 iterations) and NiO (6 vs. 2 iterations). Since each CTQMC iteration dominates the computational cost, this translates directly into a factor of 2–4 reduction in wall-clock time per structure. The learned functional thus acts as a nearly cost-free preconditioner for the DMFT loop, turning the self-energy from a repeated, expensive output into a one-shot, physics-informed initial condition.
@@
Leveraging this acceleration, we performed high-throughput DFT+DMFT calculations on an additional 600 Fe configurations at pressures and temperatures relevant to Earth’s core, increasing the Fe dataset to 1100 configurations with correlated energies and forces. Using all Fe data collected so far, we trained a NequIP-based machine-learned interatomic potential (MLIP) [nequipTan2025, nequipBatzner2022], achieving test-set mean absolute errors of 69.2 meV/atom for the energy and 76.7 meV/Å for the forces. Because the reference energies and forces are obtained from a CTQMC impurity solver, the training labels carry finite statistical uncertainty [huangMachineLearningDiffusion2023]; accordingly, the reported test errors reflect both model discrepancy and an intrinsic noise floor. In the high-temperature regime of interest (e.g., $T=5000$ K, $k_{B}T\approx 430$ meV), the resulting per-atom energy error remains sub-$k_{B}T$, enabling stable large-scale MD and solid–liquid coexistence simulations. To assess sensitivity to the electronic temperature used for DFT+DMFT labeling, we recomputed the electronic free energies for representative Fe configurations at T = 5000 and 6000 K and find changes in relative free-energy differences of $\leq$ 6 meV/atom, well below the MLIP test errors.
@@
We determined melting temperatures using the solid–liquid coexistence method [aimcAlfe2009], which avoids the hysteresis inherent in single-phase superheating or supercooling approaches. An elongated supercell (9216 atoms) containing coexisting solid (hcp) and liquid regions is equilibrated within the NVE ensemble; if the temperature lies above $T_{m}$, the solid region melts, whereas below $T_{m}$ the liquid crystallizes. The phase fractions were monitored using local bond-order analysis [pfaLechner2008]. Figure 2 illustrates a representative coexistence configuration at 307.8 GPa: the density profile along the long axis exhibits pronounced oscillations in the crystalline region, characteristic of atomic layering, and a comparatively smooth profile in the liquid region, confirming a well-defined solid–liquid interface. The temperature trace in the inset shows only moderate fluctuations after 15ps during the NVE run, indicating stable coexistence conditions. Simulations were thermalized for 8 ps and subsequently evolved for 20 ps or 30 ps, depending on phase stabilization, with a 0.5 fs timestep. Pressure was sampled from finite-difference stress calculations including the kinetic contribution, and averaged over the final 5 ps.
@@
Figure 2: Density profile along the long axis of simulation cell taken at 20 ps into equilibration for simulation at 307.8 GPa. Shown in the insets are a snapshot of the atomic configuration at 20 ps as well as the temperature variation over the length of the simulation, which stabilizes around 15 ps. The 2 coexistence interfaces can been seen on both sides of the liquid.
@@
Our melting results, spanning 307–352 GPa, is shown in Fig. 3 alongside experimental determinations and prior ab initio predictions. The results are well described by a Simon-Glatzel relation [simonglatz],
@@
with $T_{0}=6167$ K, $P_{0}=323.1$ GPa, $a=298.73$ GPa, and $c=2.43$. At the inner-core boundary pressure of 330 GPa, we obtain $T_{m}=6225\pm 42$ K, in excellent agreement with the laser-heated diamond anvil cell measurements of Anzellini et al. ($6230\pm 500$ K) [emcAnzellini2013] and the ramp-compression experiments of Kraus et al. [emcKraus2022]. It is also compatible with the recent shock-compression XAS constraints of Balugani et al., who extrapolate an upper bound of 6202 ± 514 K for Fe at 330 GPa [emcBalugani2024]. Our result lies above the resistance-heated DAC determination of Sinmyo et al. ($5500\pm 220$ K) [emcSinmyo2019]; however, as noted in Ref. [emcAnzellini2013], different melting criteria in static experiments can yield systematically different temperatures, and the discrepancy among experiments themselves remains an active area of investigation.
@@
Our agreement with recent experimental constraints also provides insight into why melting predictions for Fe at core pressures have remained so dispersed. Most published ab initio melting curves based on standard DFT span a broad range and are frequently higher than experimental estimates, with discrepancies reaching several hundred kelvin and an overall scatter approaching $10^{3}$ K [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009]. This sensitivity likely reflects, in part, that commonly used high-$T$ DFT melting workflows dynamical correlations only approximately, thereby missing high-temperature spin fluctuations that contribute to the free energy [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015].
@@
Fe at these conditions is paramagnetic, yet thermal spin fluctuations can sustain sizable instantaneous local moments and associated entropy contributions [PhysRevB.87.014405, ieccPourovskii2017, ieccVekilova2015]. If these magnetic/electronic excitations are more pronounced in the liquid than in the solid, they increase the entropy of melting and reduce the coexistence temperature through the equilibrium condition $\Delta G(P,T_{m})=0$, i.e. $T_{m}=\Delta H/\Delta S$.
@@
DFT+DMFT addresses this limitation by treating the $3d$ correlations dynamically through the frequency-dependent self-energy $\Sigma(i\omega_{n})$, capturing quasiparticle renormalization and fluctuating local moments on equal footing. That our melting curve falls within experimental uncertainty while using interaction parameters $(U,J)$ adopted from prior DFT+DMFT studies [ieccPourovskii2017, chsGavriliuk2023, chsHo2024] (rather than fitted to melting data) indicates genuine predictive power and is consistent with the view that dynamical correlations account for a substantial part of the long-standing discrepancy between simulation and experiment for Fe at core conditions.
@@
Figure 3: Full melting curve of Iron as predicted by our 2-phase simulations and compared to data from various experimental [emcAnzellini2013, emcSinmyo2019, emcLi2020, emcKraus2022, emcBalugani2024] computational studies [aimcAlfe2009, aimcGonzalez2023, aimcSun2022, aimcSun2023, aimcBelonoshko2021, aimcStixrude2014, aimcAlfe2002, aimcBelonosko2000, aimcWu2024, aimcSola2009].

### v3

# Physics-Constrained Self-Energy Warm Starts for Charge-Self-Consistent DFT+DMFT: Application to Iron at Core Conditions

@@ Abstract
Charge self-consistent DFT+DMFT quantitatively captures dynamical electronic correlations in real materials, but its cost precludes the large-scale thermodynamic sampling required for phase boundaries and equations of state. Here, we develop a physics-constrained machine-learning warm start for realistic DFT+DMFT: E(3)-equivariant graph neural networks predict a compact, real-valued representation of the local self-energy and Fermi level—{ $\Sigma(\infty),\,\Sigma_{\ell},\,E_{f}\,$}—tied to the known high-frequency and analytic structure of $\Sigma(i\omega_{n})$, and used to initialize the full DFT+DMFT self-consistency cycle. Across metallic Fe, correlated FeO, and Mott-insulating NiO, the scheme yields a 2–4 times reduction in the number of DMFT iterations required to reach self-consistency. As a demanding application, we leverage this capability to generate correlated energies and forces for Fe at core pressures, train an equivariant machine-learned interatomic potential, and determine the hcp-Fe melting curve by solid–liquid coexistence simulations in the NVE ensemble in 9216-atom cells. We obtain a melting temperature of 6225 K at 330 GPa, in agreement with recent experimental constraints and consistent with the view that dynamical electronic correlations contribute to the discrepancy between DFT-based predictions and experiment.
@@
Correlated electron systems occupy a central place in modern condensed-matter physics. Their rich behaviors, including metal–insulator transitions [1, 35], unconventional magnetism [14, 47], and high-temperature superconductivity [16, 36, 64], arise from strong local interactions, particularly in transition-metal compounds where lattice, spin, orbital, and charge degrees of freedom are tightly intertwined [4, 72]. Understanding how this interplay governs electronic properties across pressures, temperatures, and compositions remains one of the major challenges in computational materials physics and is essential for problems spanning functional oxides to materials under extreme pressure and temperature conditions.
@@
Fe at the conditions of Earth’s core exemplifies the challenges of simulating correlated systems [57, 58, 74]. The melting curve of Fe at hundreds of gigapascals controls the temperature at the inner-core boundary (ICB) and thus constrains the thermal history and dynamics of our planet, yet experimental determinations at these pressures are extremely challenging and theoretical predictions span a broad range of temperatures. While recent experiments have begun to converge on values near 6200 K at ICB pressure [5, 40, 7], direct measurement at the ICB itself remains inaccessible: laser-heated diamond-anvil cell data extrapolate upward from $\sim$200 GPa, while ramp-compression data extrapolate downward from $\sim$600 GPa, so the quoted experimental uncertainties already reflect extrapolation in addition to measurement error. The phase of Fe at melting under these conditions is also an open experimental question, with a recent preprint reporting bcc formation near melting above 200 GPa [38]. Theoretical predictions from density functional theory (DFT) calculations remain scattered, with discrepancies often exceeding 1000 K [2, 3, 10, 9, 26, 63, 65, 68, 67, 79]. A fundamental limitation of standard DFT approaches is their inadequate treatment of dynamical correlations in Fe [57, 58, 74]. Under core conditions, hexagonal close-packed (hcp) Fe and body centered cubic (bcc) Fe retain their Fermi-liquid character while fcc Fe displays markedly non-Fermi liquid behavior, with relative free energy differences changing significantly with temperature [58]. These correlation-induced differences are neglected by conventional exchange-correlation functionals but must be included for accurate simulations of liquid phases.
@@
Dynamical mean-field theory (DMFT), which captures local correlations in a fully dynamical and non-perturbative manner [39], offers a path forward. Coupled to DFT, it allows for realistic simulations of correlated materials by taking into account effects such as quasiparticle renormalizations, and provides access to electronic and thermodynamic properties such as spectral functions and finite-temperature phase diagrams [56]. However, this power comes at a substantial computational cost. Each DMFT calculation requires repeated solution of the quantum impurity problem until impurity and lattice occupations become consistent. Determining a melting curve requires sampling thousands of large-scale atomic configurations through molecular dynamics, a task that has remained prohibitively expensive within DFT+DMFT.
@@
Recently, machine learning has begun to be explored as a way to reduce the cost of DMFT calculations, for example through surrogate impurity solvers, learned corrections to approximate Green’s functions, and data-driven models trained on simplified parameter spaces [51, 20, 66, 6, 73, 53]. While these studies demonstrate that key DMFT quantities can often be approximated efficiently in selected settings, most existing approaches have been confined to model Hamiltonians or narrow material classes, and the learned targets are not always constructed to respect the analytic structure and symmetry constraints required in realistic DFT+DMFT workflows. It therefore remains unclear whether machine learning can provide a robust, physics-consistent acceleration of DFT+DMFT for correlated solids with the accuracy needed for thermodynamic predictions, particularly under the extreme conditions relevant to Earth’s core.
@@
In this work, we demonstrate that embedding physical constraints into machine-learned self-energies enables scalable finite-temperature DFT+DMFT thermodynamics for correlated materials, and we use this capability to determine the hcp-Fe melting curve at core conditions as a demanding application. Our approach exploits the analytic structure of the self-energy: we decompose it into a static high-frequency limit $\Sigma(\infty)$ and a frequency-dependent term $\Sigma(i\omega)$, represented compactly through Legendre coefficients, with each component predicted by E(3)-equivariant graph neural networks that respect the point-group symmetry of the local atomic environment. This physics-informed representation reduces the learning problem to smooth, real-valued targets that respect the analytic structure required by the DFT+DMFT self-consistency cycle [6].
@@
The resulting model accelerates DMFT convergence by a factor of two to four across Fe, FeO, and NiO, enabling high-throughput generation of correlated electronic structure data. We leverage this capability to train a machine learned interatomic potential (MLIP) for iron and perform solid–liquid coexistence simulations in the NVE ensemble [52, 3] spanning 307–352 GPa. Our predicted melting temperature of 6225K at 330 GPa agrees well with the recent experimental measurements [5, 40, 7], providing strong validation that large-scale molecular dynamics (MD) simulations using DMFT-based MLIPs are viable for phase-equilibrium calculations in planetary interiors.
@@
A naive strategy would be to train a neural network to output $\Sigma(i\omega_{n})$ directly on a discrete Matsubara frequency grid. This is undesirable for several reasons. First, the DMFT self-energy is complex-valued and, when obtained from quantum Monte Carlo impurity solvers, can be noisy at high frequencies. Second, $\Sigma(i\omega_{n})$ obeys strict analytic constraints: it is a causal function with a well-defined high-frequency expansion, $\Sigma(i\omega_{n})=\Sigma(\infty)+{\cal O}(1/i\omega_{n})$ [76], and its imaginary part has definite sign on the imaginary axis. Direct regression on a large set of complex values risks violating these constraints unless they are enforced explicitly.
@@
[…]
@@
The dynamical part $\Delta\Sigma(i\omega)$ is then Fourier transformed to imaginary time and expanded in a Legendre polynomial basis [13],
@@
where $\Delta\Sigma^{\mathrm{ML}}(i\omega_{n})$ is obtained by transforming the Legendre expansion back to the Matsubara axis. Both forward and reverse transformations are performed by the TRIQS software package [55, 13]. This $\Sigma^{\mathrm{ML}}$, together with the predicted $E_{f}$, provides the initial condition for the DMFT self-consistency loop.
Including $E_{f}$ as a learned quantity is essential for achieving substantial speedups. In DFT+DMFT the Fermi level is normally adjusted iteratively so that the total electron number satisfies the charge constraint, and poor initial guesses can severely delay convergence. By treating $E_{f}$ on the same footing as $\Sigma(\infty)$ and the Legendre coefficients, we define a joint functional $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ that maps a crystal structure to a complete, physics-consistent set of input parameters.
To implement this functional, we represent each crystal structure as a periodic atomic graph and approximate $\left\{\Sigma^{\mathrm{ML}}[{\mathcal{S}}],E_{f}^{\mathrm{ML}}[{\mathcal{S}}]\right\}$ with an $E(3)$-equivariant graph neural network built using the open-source e3nn software package [25, 37, 71, 77]. Atoms are nodes, connected to neighbors within a radial cutoff, and are encoded by learned embeddings of their chemical species and mass. Relative position vectors between neighbors are expanded in spherical harmonics and radial basis functions, so that all edge features and intermediate representations transform as irreducible representations of $E(3)$. A small stack of equivariant message-passing layers then builds a symmetry-respecting description of each local environment that is translationally invariant and equivariant under rotations and inversion. Further architectural details are given in the Supplemental Material.
@@
To assess the accuracy of the physics-informed functional, we first compare its predictions for the self-energy against fully converged DFT+DMFT results on held-out test structures. For Fe and FeO, structural configurations were generated from MD simulations of high-symmetry crystalline prototypes (e.g., hcp/bcc/fcc for Fe, B2 phase for FeO) to ensure coverage of both ordered and disordered environments. 2.5-ps DFT-MD trajectories (0.5 fs timestep) of 4–8 atom cells were generated and representative snapshots were selected by k-medoids clustering in structural fingerprint space [80], yielding approximately 500 configurations per compound (100–200 per seed trajectory). Since the fingerprint represents the chemical environment of each crystal structure, the fingerprint distance allows us to quantify the overall difference in chemical environments between any two crystal structures. This ensures that we obtain diverse chemical environments for the solver to “learn” in an attempt to aid generalization of results to unseen structures. For NiO we use configurations from our previous study [60]. Full details of dataset generation and selection are provided in the Supplemental Material. We reserve 10% as a held-out test set.
@@
Figure 1 (d-f) shows the convergence of impurity-lattice occupation difference $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|$ as a function of DMFT iteration for Fe, FeO, and NiO, averaged over the data for 20 randomly chosen structures. With zero initialization, Fe requires approximately 9 iterations to reach our convergence criterion of $|n_{\mathrm{imp}}-n_{\mathrm{lat}}|<0.05$; with ML-predicted self-energies, convergence is achieved in 2–3 iterations. Similar speedups are observed for FeO (9 vs. 5 iterations) and NiO (6 vs. 2 iterations). Since CTQMC iterations dominate the computational cost, this translates directly into a factor of 2–4 reduction in wall-clock time per structure. The learned functional thus acts as a nearly cost-free preconditioner for the DMFT loop, turning the self-energy from a repeated, expensive output into a one-shot, physics-informed initial condition.
FeO retains the largest iteration count under ML initialization, reflecting the complex Fe 3d3d–O $2p$ hybridization in the B2 phase and the correspondingly stronger frequency dependence of the self-energy, which makes the DMFT loop more sensitive to residual errors in the higher-order Legendre coefficients (see Supplemental Material).
@@
Leveraging this acceleration, we performed high-throughput DFT+DMFT calculations on an additional 600 Fe configurations at pressures and temperatures relevant to Earth’s core, increasing the Fe dataset to 1100 configurations with correlated energies and forces. Using all Fe data collected so far, we trained a NequIP-based machine-learned interatomic potential (MLIP) [69, 8], achieving test-set mean absolute errors of 69.2 meV/atom for the energy and 76.7 meV/Å for the forces. Because the reference energies and forces are obtained from a CTQMC impurity solver, the training labels carry finite statistical uncertainty [34]; accordingly, the reported test errors reflect both model discrepancy and an intrinsic noise floor. In the high-temperature regime of interest (e.g., $T=5000$ K, $k_{B}T\approx 430$ meV), the resulting per-atom energy error remains sub-$k_{B}T$, enabling stable large-scale MD and solid–liquid coexistence simulations.
The per-atom global energy MAE does not directly determine the melting-temperature error; the relevant quantity is the phase-distinguishing signed residual in the solid–liquid energy/free-energy difference. A phase-resolved analysis on the held-out Fe test set gives a central $\delta T_{m}$ estimate of $\sim 20$ K, while alternative phase partitions give a conservative systematic envelope of order $10^{2}$ K, reported separately from the $\pm 42$ K internal coexistence fluctuation (see Supplemental Material for the full propagation, including the $L_{\mathrm{melt}}=T_{m}\,\Delta S_{m}$ derivation [2, 3]). The MLIP force MAE of 76.7 meV/Å is comparable to or lower than values reported in recent finite-temperature and extreme-condition MLIP studies, including melting and crystallization applications [18, 17, 22, 59, 70, 78], supporting the fidelity of the coexistence trajectories. Residual stress anisotropy of 2–3 GPa from the fixed-cell NVE protocol contributes a separate pressure-axis uncertainty (see Supplemental Material). To assess sensitivity to the labelling temperature we recomputed representative relative free-energy differences at 5000 and 6000 K and find changes of $\leq 6$ meV/atom, well below the MLIP test errors.
@@
We determined melting temperatures using solid–liquid coexistence simulations in the NVE ensemble [52, 3], which avoid the hysteresis inherent in single-phase superheating or supercooling approaches. An elongated supercell (9216 atoms) containing coexisting solid (hcp) and liquid regions is equilibrated within the NVE ensemble; if the temperature lies above $T_{m}$, the solid region melts, whereas below $T_{m}$ the liquid crystallizes. The phase fractions were monitored using local bond-order analysis [46]. Figure 2 illustrates a representative coexistence configuration at 307.8 GPa: the density profile along the long axis exhibits pronounced oscillations in the crystalline region, characteristic of atomic layering, and a comparatively smooth profile in the liquid region, confirming a well-defined solid–liquid interface. The temperature trace in the inset shows only moderate fluctuations after 15ps during the NVE run, indicating stable coexistence conditions. Simulations were thermalized for 8 ps and subsequently evolved for 20 ps or 30 ps, depending on phase stabilization, with a 0.5 fs timestep. Pressure was sampled from finite-difference stress calculations including the kinetic contribution, and averaged over the final 5 ps. To verify that the MLIP remains well-behaved in the disordered interface region, we rely on two diagnostics: the training set already contains liquid-like and high-temperature solid configurations sampled from DFT-MD seeds run at or above the small-cell melting temperature, and the coexistence MD itself shows stable behaviour—over tens of picoseconds the interface remains stationary, the NVE energy drift is consistent with integration error rather than with model instability, and the interfacial density profile is smooth (see Supplemental Material).
@@
Figure 2: Density profile along the long axis of simulation cell taken at 20 ps into equilibration for simulation at 307.8 GPa. Shown in the insets are a snapshot of the atomic configuration at 20 ps as well as the temperature variation over the length of the simulation, which stabilizes around 15 ps. The 2 coexistence interfaces can be seen on both sides of the liquid.
@@
Our melting results, spanning 307–352 GPa, are shown in Fig. 3 alongside experimental determinations and prior ab initio predictions. The results are well described by a Simon-Glatzel relation [11],
@@
with $T_{0}=6167$ K, $P_{0}=323.1$ GPa, $a=298.73$ GPa, and $c=2.43$. At the inner-core boundary pressure of 330 GPa, we obtain $T_{m}=6225\pm 42\,\mathrm{K}_{\mathrm{stat}}\pm\mathcal{O}(10^{2})\,\mathrm{K}_{\mathrm{sys}}$, where the first bar reflects the internal fluctuation of the coexistence MD over the final 5 ps and the second is the systematic uncertainty from the phase-distinguishing component of the MLIP error described above (Supplemental Material). This is in agreement with the laser-heated diamond anvil cell measurements of Anzellini et al. ($6230\pm 500$ K) [5] and the ramp-compression experiments of Kraus et al. [40], both of which rely on extrapolation to the ICB (upward from $\sim$200 GPa and downward from $\sim$600 GPa, respectively), so that the quoted experimental bars themselves already include substantial extrapolation uncertainty. Our result is also compatible with the recent shock-compression XAS constraints of Balugani et al., who extrapolate an upper bound of 6202 $\pm$ 514 K for Fe at 330 GPa [7]. It lies above the resistance-heated DAC determination of Sinmyo et al. ($5500\pm 220$ K) [62]; however, as noted in Ref. [5], different melting criteria in static experiments can yield systematically different temperatures, and the discrepancy among experiments themselves remains an active area of investigation.
@@
Our agreement with recent experimental constraints also provides insight into why melting predictions for Fe at core pressures have remained so dispersed. Most published ab initio melting curves based on standard DFT span a broad range and are frequently higher than experimental estimates, with discrepancies reaching several hundred kelvin and an overall scatter approaching $10^{3}$ K [3, 26, 68, 67, 10, 65, 2, 9, 79, 63]. This sensitivity likely reflects, in part, that commonly used high-$T$ DFT melting workflows incorporate dynamical correlations only approximately, thereby missing high-temperature spin fluctuations that contribute to the free energy [61, 57, 74].
@@
Fe at these conditions is paramagnetic, yet thermal spin fluctuations can sustain sizable instantaneous local moments and associated entropy contributions [61, 57, 74]. If these magnetic/electronic excitations are more pronounced in the liquid than in the solid, they increase the entropy of melting and reduce the coexistence temperature through the equilibrium condition $\Delta G(P,T_{m})=0$, i.e. $T_{m}=\Delta H/\Delta S$.
@@
DFT+DMFT addresses this limitation by treating the $3d$ correlations dynamically through the frequency-dependent self-energy $\Sigma(i\omega_{n})$, capturing quasiparticle renormalization and fluctuating local moments on equal footing. That our melting curve falls within experimental uncertainty while using interaction parameters $(U,J)$ adopted from prior DFT+DMFT studies [57, 23, 33] (rather than fitted to melting data) indicates genuine predictive power and is consistent with the view that dynamical correlations account for a substantial part of the long-standing discrepancy between simulation and experiment for Fe at core conditions.
@@
Figure 3: Full melting curve of Iron as predicted by our solid–liquid coexistence simulations and compared to data from various experimental [5, 62, 48, 40, 7] computational studies [3, 26, 68, 67, 10, 65, 2, 9, 79, 63].